Calculate the standard deviation for Stock C. (Enter percentages as decimals and round to 4 decimals)

State Prob(State) Stock A Stock B Stock C
Boom 30% -12% 16% -3%
Modest Growth 50% 30% 5% -1%
Recession 20% -12% -1% 25%
Answer #1

Average Return=Respective Return*Respective probability

=(0.3*-3)+(0.5*-1)+(0.2*25)=3.6%

probability Return probability*(Return-Expected Return)^2
0.3 -3 0.3*(-3-3.6)^2=13.068
0.5 -1 0.5*(-1-3.6)^2=10.58
0.2 25 0.2*(25-3.6)^2=91.592
Total=115.24%

Standard deviation=[Total probability*(Return-Expected Return)^2/Total probability]^(1/2)
=0.1073(Approx).